OpenStudy (anonymous):

Discount tickets to a basketbal tournament sell for $4.00 each. Enrico spent $60.00 on discount tickets, $37.50 less than if he had bought the tickets at the regular price. What was the regular ticket price?

OpenStudy (anonymous):

First determine he bought 15 tickets...with discount total (60/4) 60 + 37.50 = 97.50/15 = 6.50 (regular price)

OpenStudy (anonymous):

discounted price of one ticket=$4 no. of tickets=60/4=15